# A Dynamic, Self-balancing k-d Tree

### v13

@@ 2 Conclusions
The dynamic k-d tree is implemented as the KdTreeDynamic class that derives from (in C++), or extends (in Java), the KdTree class that implements the $O\left[n\log\left(n\right)\right]$ and $O\left[kn\log\left(n\right)\right]$ algorithms that build a static k-d tree. Hence, all other algorithms implemented by the KdTree class, such as region search, nearest-neighbor search and reverse-nearest-neighbor search [Korn], are accessible by the KdTreeDynamic class. In addition, it is possible to build a static k-d tree via an $O\left[n\log\left(n\right)\right]$ or $O\left[kn\log\left(n\right)\right]$ algorithm, and then provide that static tree to the KdTreeDynamic constructor so that the static tree may thereafter be updated dynamically via insertion or deletion of individual tuples.

### v14

@@ 2 Conclusions
Future research could compare the average height to the maximum height at each node of k-d trees that are balanced via the AVL or red-black criteria to determine whether the average height correlates with the maximum height.
@@
The dynamic k-d tree is implemented as the KdTreeDynamic class that derives from (in C++), or extends (in Java), the KdTree class that implements the $O\left[n\log\left(n\right)\right]$ and $O\left[kn\log\left(n\right)\right]$ algorithms that build a static k-d tree. Hence, all other algorithms implemented by the KdTree class, such as region search, nearest-neighbor search and reverse-nearest-neighbor search [10], are accessible by the KdTreeDynamic class. In addition, it is possible to build a static k-d tree via an $O\left[n\log\left(n\right)\right]$ or $O\left[kn\log\left(n\right)\right]$ algorithm, and then provide that static tree to the KdTreeDynamic constructor so that the static tree may thereafter be updated dynamically via insertion or deletion of individual tuples.